Two Particles A & B are projected from the same point with the same velocity u of projection but at different angles (x) & (y), such that the maximum height of A is two-third of the horizontal range of B.Then which of the following relations are true ?


(a) Range of A = Maximum height of B


(b) Maximum horizontal range of A = u^2/g and this occurs when (y) = 1/2 sin^-1 3/8


(c) 3(1 - cos 2{x}) = 8 sin 2(y)


(d) Maximum value of (y) is 1/2 sin^-1 3/4


 


[Ans: (b) , (c) , (d)]


 


Grade:11

1 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

Dear   zahid sharief,

The initial velocity v0

v_0=v_{0x}\mathbf{i} + v_{0y}\mathbf{j}

 if the angle θ0 is known:

v0x = v0cosθ0
v0y = v0sinθ0

 

 

The horizontal displacement(xx0 )

 

x-x_0=v_{0x}t+\begin{matrix}\frac{1}{2}\end{matrix}at^2

 

xx0 = (v0cosθ0)t.

.y-y_0=(v_0\sin\theta_0)t-\begin{matrix}\frac{1}{2}\end{matrix}gt^2.

 

y-y_0=(v_0\sin\theta_0)t-\begin{matrix}\frac{1}{2}\end{matrix}gt^2


equation of projectile

y=x\tan(\theta)-x^2g/(2v^2_{0}cos^2(\theta))

Time to reach the maximum height

t={v_{0y}\over g}

Range of projectile

{r=\frac{v^2}{g}\sin(2\theta)}\,

Maximum height                                         

y_{max}={{v^2\sin^2\theta}\over 2g}

 

 

maimum height of  A     y1(max) =   v^2 (1-cos 2x)/4 *g       here  2 sin2x= 1-cos2x

range of                        r2(max) = v^2*sin(2y)/g     

 

compare them we get    3(1 - cos 2{x}) = 8 sin 2(y)

 

maimum value    (y)= 1/2 [sin-1 {3/8(1 - cos 2{x}) } ]

 

say cos2x=-1

we get maximum value   that is     max(y)= 1/2 [sin-1 {3*2/8} ]

                                   (y)= 1/2 [sin-1 3/4 ]

 

 

 


Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar..



Askiitians Expert

Rathod Shankar

IIT bombay